7. At 20°C, a 2.32 M aqueous solution of ammonium chloride has a density of 1.0344 g/mL. What is the molality of ammonium chloride in the solution? The formula weightof NH4Cl is 53.50 g/mol
Answer #1
Step1 Mass of 1 litre of solution = 1000x1.0344 = 1034.4 g

Step2 Mass of solute in 1 litre of solution = 2.32x 53.5=124.12 g

Step3 Mass of solvent= 1034.4-124.12=910.28

Step4 Molality = Moles per kg solvent = (2.32)x1000/910.28 =2.55